# Active viscoelastic condensates provide controllable mechanical anchor points

### v1

@@ Conclusions
The precise and controlled assembly of mechanical anchor points is vital to the function of a range of biological and synthetic materials, including biomolecular condensates. In this manuscript, we have developed a continuum model that demonstrates how viscoelastic material properties, coupled with actively driven incorporation of new material around a central core, can facilitate the formation of such structures. Our model thus provides a theoretical framework with which to analyse the growth and deformation of these objects, and to constrain the material properties that are consistent with experimental observations. Although viscoelastic stresses restrict condensate growth, we find that it is possible to combine both rapid growth and mechanical strength in an appropriate parameter regime. We also demonstrate that localised incorporation of new condensate material in the vicinity of the active core may be suppressed if incorporation rates depend on condensate stress state. Finally, we have demonstrated the applicability of our theory by comparing our model predictions to previous experimental data for C. elegans centrosomes, identifying allowable values of different material parameters. Taken together, our findings suggest that viscoelastic condensates provide a simple solution for cells to rapidly form mechanical anchor points.
@@
In this work, we have made several simplifying approximations to focus on the core physical principles that govern the formation, growth and deformation of viscoelastic condensates in cells. Relaxing these assumptions will facilitate detailed investigation of the effects of physical phenomena not considered here. For example, we have focussed primarily on the limiting case in which precursor material distributes rapidly and uniformly throughout the cell. Our integrated model can be straightforwardly extended to situations in which the finite diffusion time of precursor is important by explicitly tracking the diffusion-governed motion of precursor material, as we already do in our phase-field simulations. Further, the influence of time-dependent material properties [e.g., 31, 32], could be trivially incorporated into our model given a functional dependence of material stiffness or relaxation time on condensate age. Lastly, although our integrated model neglects variations in the spatial distribution of scaffold material inside a condensate, the impact of gradients in scaffold deformation (e.g., due to poroelastic drag) can be readily explored via the detailed phase-field simulations presented in figure 2. As such, we expect that the general framework developed here will be useful for studying a wide range of systems, including other examples of biomolecular condensates (e.g., focal adhesions and tight junctions), as well as tissues at the multi-cellular scale, and biomimetic materials such as bio-adhesives and tissue engineering scaffolds [33].
@@ Acknowledgements
The authors thank Sebastian Fürthauer for helpful discussions. OWP and DZ acknowledge funding from the Max Planck Society and the European Union (ERC, EmulSim, 101044662). This work was also supported by funding from the Austrian Science Fund FWF, grant P34526-B, to AD, as well as a Max Perutz PhD fellowship of the Max Perutz Labs to JG-B. SA and LZ acknowledge support from DFG (grant 511509575). SA and LZ also gratefully acknowledge computing time on the high-performance computer at the NHR Center of TU Dresden. This center is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner). This research was supported in part by grant NSF PHY-2309135 and the Gordon and Betty Moore Foundation Grant No. 2919.02 to the Kavli Institute for Theoretical Physics (KITP).
@@ A.1 Phase-field formulation of condensate growth
$K_{\mathrm{b}}(\tilde{\phi})$ $=K_{\mathrm{b}}\tilde{\phi}\;,$
$K_{\mathrm{s}}(\tilde{\phi})$ $=K_{\mathrm{s}}\tilde{\phi}\;.$
@@ A.2 Summary of spherically symmetric equations
[…]
@@
$\partial_{t}\phi_{P}$ $=m_{P}\partial_{r}\left(\phi_{P}\partial_{r}\mu_{P}\right)+\frac{2}{r}m_{P}\phi_{P}\partial_{r}\mu_{P}-s\;,$ (18)
$\partial_{t}\phi_{S}$ $=m_{S}\partial_{r}\left(\phi_{S}\partial_{r}\mu_{S}\right)+\frac{2}{r}m_{S}\phi_{S}\partial_{r}\mu_{S}-Em_{S}\partial_{r}\left(\partial_{r}\sigma_{rr}+\frac{2}{r}(\sigma_{rr}-\sigma_{\theta\theta})\right)+s\;,$ (19)
@@ A.3 Details of numerical method
${C_{h}}_{1}$ =$=\left\{c\in C^{0}(\bar{\Omega})\mid c|_{k}\in P_{2}(k),k\in T_{\Omega}\right\}\;,$
${C_{h}}_{2}$ =$=\left\{c\in C^{0}(\bar{\Omega})\mid c|_{k}\in P_{1}(k),k\in T_{\Omega}\right\}\;,$
$C_{q,n}$ =$=\left\{c\in C^{0}(\bar{\Omega})\mid c|_{k}\in P_{1}(k),k\in T_{\Omega},c(r_{\textrm{c}})=q\phi_{P}^{n-1}(r_{\textrm{c}})/\phi_{S}^{n-1}(r_{\textrm{c}}),c(r_{\textrm{d}})=0\right\}$
@@
$\int_{\Omega}\frac{\phi_{P}^{n}-\phi_{P}^{n-1}}{\Delta t}w_{1}\mathop{}\!\mathrm{d}r$ $=-m_{P}\int_{\Omega}\phi_{P}^{n-1}\partial_{r}\mu_{P}^{n}\partial_{r}w_{1}\mathop{}\!\mathrm{d}r+q\phi_{P}^{n-1}(r_{c})w_{1}(r_{c})+m_{P}\int_{\Omega}\frac{2}{r}\phi_{P}^{n-1}\partial_{r}\mu_{P}^{n}w_{1}\mathop{}\!\mathrm{d}r$
$\hskip 10.00002pt-k_{+}\int_{\Omega}\tilde{\phi}^{n-1}\phi_{P}^{n}w_{1}\mathop{}\!\mathrm{d}r+k_{-}\int_{\Omega}\phi_{S}^{n}w_{1}\mathop{}\!\mathrm{d}r,\;$
@@
$\int_{\Omega}\frac{\phi_{S}^{n}-\phi_{S}^{n-1}}{\Delta t}w_{3}\mathop{}\!\mathrm{d}r$ $=-m_{S}\int_{\Omega}\phi_{S}^{n-1}\partial_{r}\mu_{S}^{n}\partial_{r}w_{3}\mathop{}\!\mathrm{d}r-q\phi_{P}^{n-1}(r_{c})w_{1}(r_{c})+m_{S}\int_{\Omega}\frac{2}{r}\phi_{S}^{n-1}\partial_{r}\mu_{S}^{n}w_{3}\mathop{}\!\mathrm{d}r$
$\hskip 10.00002pt+Em_{S}\int_{\Omega}\left(\partial_{r}\sigma_{rr}^{n}+\frac{2}{r}(\sigma_{rr}^{n}-\sigma_{\theta\theta}^{n})\right)\partial_{r}w_{3}\mathop{}\!\mathrm{d}r+k_{+}\int_{\Omega}\tilde{\phi}^{n-1}\phi_{P}^{n}w_{3}\mathop{}\!\mathrm{d}r-k_{-}\int_{\Omega}\phi_{S}^{n}w_{3}\mathop{}\!\mathrm{d}r\;,$
@@
The simulation parameters used for the phase-field results presented in figure 2 are provided in table 1.
@@
Domain $\Omega$ $[r_{\textrm{c}},r_{\textrm{d}}]$
@@
[…]
@@
[…]
@@
[…]
@@ B.1 Detailed derivation of integrated model
$-\mathbf{T}\cdot\hat{\bm{r}}=\sigma^{*}$ (35)

### v2

@@ IV Conclusions
The precise and controlled assembly of mechanical anchor points is vital to the function of a range of biological and synthetic materials, including biomolecular condensates. In this manuscript, we have developed a continuum model that demonstrates how viscoelastic material properties, coupled with actively driven incorporation of new material around a central core, can facilitate the formation of such structures. Our model thus provides a theoretical framework with which to analyse the growth and deformation of these objects, and to constrain the material properties that are consistent with experimental observations. Although viscoelastic stresses restrict condensate growth, we find that it is possible to combine both rapid growth and mechanical strength in an appropriate parameter regime. We also demonstrate that localised incorporation of new condensate material in the vicinity of the active core may be suppressed if incorporation rates depend on condensate stress state. Finally, we have demonstrated the applicability of our theory by comparing our model predictions to experimental data for C. elegans centrosomes [28], quantitatively capturing the growth dynamics of both wild-type and nocodazole-treated samples. Taken together, our findings suggest that active viscoelastic condensates provide a simple solution for cells to rapidly form mechanical anchor points.
The application of our modelling framework to the case of centrosomes represents a significant development on previous theories of centrosome growth [e.g. 17, 31], reconciling the fact that centrosomes must display suitable mechanical strength in addition to rapid maturation, and identifying a parameter regime in which this is possible. The broad range of allowable elastic moduli suggests that this mechanism is robust even allowing for recently observed softening of the PCM in the approach to mitosis [28]. In addition, we have shown that elastic stresses may inhibit the incorporation of new scaffold material, particularly in the vicinity of the centriole, where stresses are largest. This effect could counteract the expected enhanced incorporation resulting from a diffusible signal emanating from the centriole [32], potentially explaining previously observed isotropic incorporation [33]. In summary, our model can explain many aspects of centrosome formation and growth, and lays a key foundation for developing future understanding.
@@
In this work, we have made several simplifying approximations to focus on the core physical principles that govern the formation, growth and deformation of viscoelastic condensates in cells. Relaxing these assumptions will facilitate detailed investigation of the effects of physical phenomena not considered here. For example, we have focussed primarily on the limiting case in which precursor material distributes rapidly and uniformly throughout the cell. Our integrated model can be straightforwardly extended to situations in which the finite diffusion time of precursor is important by explicitly tracking the diffusion-governed motion of precursor material, as we already do in our phase-field simulations. Further, the influence of time-dependent material properties [e.g., 30, 34, 35, 28], could be trivially incorporated into our model given a functional dependence of material stiffness or relaxation time on condensate age. Lastly, although our integrated model neglects variations in the spatial distribution of scaffold material inside a condensate, the impact of gradients in scaffold deformation (e.g., due to poroelastic drag) can be readily explored via the detailed phase-field simulations presented in figure 2. As such, we expect that the general framework developed here will be useful for studying a wide range of systems, including other examples of biomolecular condensates (e.g., focal adhesions and tight junctions), as well as tissues at the multi-cellular scale, and biomimetic materials such as bio-adhesives and tissue engineering scaffolds [36].
@@ Acknowledgements
The authors thank Sebastian Fürthauer for helpful discussions. OWP and DZ acknowledge funding from the Max Planck Society and the European Union (ERC, EmulSim, 101044662). This work was also supported by funding from the Austrian Science Fund FWF, grant P34526-B, to AD, as well as a Max Perutz PhD fellowship of the Max Perutz Labs to JG-B. SA and LZ acknowledge support from DFG (grant 511509575). SA and LZ also gratefully acknowledge computing time on the high-performance computer at the NHR Center of TU Dresden. This centre is jointly supported by the Federal Ministry of Education and Research and the state governments participating in the NHR (www.nhr-verein.de/unsere-partner). This research was supported in part by grant NSF PHY-2309135 and the Gordon and Betty Moore Foundation Grant No. 2919.02 to the Kavli Institute for Theoretical Physics (KITP).
Author contributions
OWP and DZ designed research; OWP, LZ, SA, and DZ developed the model; OWP and LZ performed simulations; JG-B and AD designed and performed experiments; OWP analysed data; OWP wrote the original draft; and all authors edited the final manuscript.
@@ A.1 Phase-field formulation of condensate growth
[…]
@@ A.2 Summary of spherically symmetric equations
$\begin{rcases}-\partial_{r}\mu_{P}&=-\frac{q}{m_{P}}\\ -\partial_{r}\mu_{S}&=\frac{q\phi_{P}}{m_{S}\phi_{S}}\\ v&=q\phi_{P}/\phi_{S}\\ \partial_{r}B_{rr}&=0\\ \partial_{r}B_{\theta\theta}&=0\end{rcases}$ $\text{ at }r=r_{\textrm{c}}\;.$ (16)
@@
$r^{2}\partial_{t}\phi_{P}$ $=m_{P}\partial_{r}\left(r^{2}\phi_{P}\partial_{r}\mu_{P}\right)-r^{2}s\;,$ (18)
$r^{2}\partial_{t}\phi_{S}$ $=m_{S}\partial_{r}\left(r^{2}\phi_{S}\partial_{r}\mu_{S}\right)-Em_{S}\partial_{r}\left(r^{2}\left(\partial_{r}\sigma_{rr}+\frac{2}{r}(\sigma_{rr}-\sigma_{\theta\theta})\right)\right)+r^{2}s\;,$ (19)
@@ A.3 Details of numerical method
${C_{h}}_{1}$ $=\left\{c\in C^{0}(\bar{\Omega})\mid c|_{k}\in P_{2}(k),k\in T_{\Omega}\right\}\;,$
${C_{h}}_{2}$ $=\left\{c\in C^{0}(\bar{\Omega})\mid c|_{k}\in P_{1}(k),k\in T_{\Omega}\right\}\;,$
$C_{q,n}$ $=\left\{c\in C^{0}(\bar{\Omega})\mid c|_{k}\in P_{1}(k),k\in T_{\Omega},c(r_{\textrm{c}})=q\phi_{P}^{n-1}(r_{\textrm{c}})/\phi_{S}^{n-1}(r_{\textrm{c}}),c(r_{\textrm{d}})=0\right\}$
@@
$\int_{\Omega}r^{2}\frac{\phi_{P}^{n}-\phi_{P}^{n-1}}{\Delta t}w_{1}\mathop{}\!\mathrm{d}r$ $=-m_{P}\int_{\Omega}r^{2}\phi_{P}^{n-1}\partial_{r}\mu_{P}^{n}\partial_{r}w_{1}\mathop{}\!\mathrm{d}r+qr^{2}\phi_{P}^{n-1}(r_{c})w_{1}(r_{c})$
$\hskip 10.00002pt-k_{+}\int_{\Omega}r^{2}\tilde{\phi}^{n-1}\phi_{P}^{n}w_{1}\mathop{}\!\mathrm{d}r+k_{-}\int_{\Omega}r^{2}\phi_{S}^{n}w_{1}\mathop{}\!\mathrm{d}r,\;$
@@
$\int_{\Omega}r^{2}\frac{\phi_{S}^{n}-\phi_{S}^{n-1}}{\Delta t}w_{3}\mathop{}\!\mathrm{d}r$ $=-m_{S}\int_{\Omega}r^{2}\phi_{S}^{n-1}\partial_{r}\mu_{S}^{n}\partial_{r}w_{3}\mathop{}\!\mathrm{d}r-qr^{2}\phi_{P}^{n-1}(r_{c})w_{1}(r_{c})$
$\hskip 10.00002pt+Em_{S}\int_{\Omega}r^{2}\left(\partial_{r}\sigma_{rr}^{n}+\frac{2}{r}(\sigma_{rr}^{n}-\sigma_{\theta\theta}^{n})\right)\partial_{r}w_{3}\mathop{}\!\mathrm{d}r+k_{+}\int_{\Omega}r^{2}\tilde{\phi}^{n-1}\phi_{P}^{n}w_{3}\mathop{}\!\mathrm{d}r-k_{-}\int_{\Omega}r^{2}\phi_{S}^{n}w_{3}\mathop{}\!\mathrm{d}r\;,$
@@
The simulation parameters used for the phase-field results presented in Fig. 2 are provided in table  1.
@@
Domain $\Omega$ $r_{\textrm{c}},r_{\textrm{d}}]$
@@
Time step size $\Delta t$ $0.0025\text{\,}\mathrm{s}$
@@
Rate of conversion of scaffold into precursor $k_{-}$ $0.001\text{\,}{\mathrm{s}}^{-1}$
Rate of conversion of precursor into scaffold in the bulk of the condensate $k_{+}$ $0.4\text{\,}{\mathrm{s}}^{-1}$
Rate of conversion of precursor into scaffold at the surface of the active core $q$ $4.2\text{\times}{10}^{-8}\text{\,}{\mathrm{s}}^{-1}$
@@
Bulk modulus of the scaffold $K_{\mathrm{b}}$ $1\text{\,}\mathrm{N}\text{\,}{\mathrm{m}}^{-2}$
Shear modulus of the scaffold $K_{\mathrm{s}}$ $1\text{\,}\mathrm{N}\text{\,}{\mathrm{m}}^{-2}$
Relaxation time in the bulk of the condensate $\tau_{-}$ $100\text{\,}\mathrm{s}$
Relaxation time in the surrounding environment $\tau_{+}$ $0.1\text{\,}\mathrm{s}$
@@ B.1 Detailed derivation of integrated model
$\mathbf{T}\cdot\hat{\bm{r}}=\sigma^{*}$ (35)
@@ B.3 Deformation of a viscoelastic sphere
Appendix C Experimental details
The quantitative analysis of SPD-5 PCM scaffold polymer expansion in one-, two- and four-cell stage C. elegans embryos presented in Fig. 6 is based on experimental data described in the accompanying manuscript by Garcia-Baucells et al. [28]. We briefly outline the experimental methods below. For further details on the methods used please refer to that manuscript.
C.1 Materials and methods
Experiments were performed in worms co-expressing GFP:SPD-5, mCherry:SPD-2 and mCherry:Histone (strain DAM946: spd-5(vie26[gfp::spd-5 +loxP]) I; unc-119(ed3) III; ltIs37[pAA64; Ppie-1::mcherry::his-58; unc-119(+)]; ltIs69[pAA191; Ppie-1::mcherry::spd-2; unc-119(+)] IV [32]. To enable nocodazole-treatment of early embryos, egg shells were permeabilised using perm-1 RNAi [40], while simultaneously inhibiting the spindle checkpoint with $\textrm{mad-}2^{\textrm{mdf-2}}$} RNAi to prevent mitotic arrest [41]. RNAi was performed by soaking [42], with a recovery time of 24 hours at 20 °C to minimise phenotypes beyond embryo permeability.
C.2 Imaging
[…]
C.3 Nocodazole treatment
[…]
C.4 Image processing
Post image acquisition, image stacks were processed using Fiji/ImageJ (version 1.54f), separating channels and extracting timestamp metadata, with subsequent analysis performed in MATLAB (version R2021b). Movies were aligned with respect to two reference time points: (1) male pronuclear/nuclear permeabilisation for S-phase-treated centrosomes, and (2) chromosome decondensation for NEBD-treated embryos. Nuclear permeabilisation was defined as the last time point preceding complete cytoplasmic GFP invasion of the nuclear space. Chromosome decondensation timing was determined using a modified version of the method by Maddox et al. [43]. Prior to analysis, images were corrected for spherical aberration, camera noise and photobleaching.
[…]